OpenStudy (anonymous):

Factor the polynomial: x^(2) - 121

OpenStudy (anonymous):

Also: t^(3) + 2t^(2) - 9t - 18

OpenStudy (anonymous):

121 is nothing but \(11^2\).. So now you can write it as: \(x^2 - 11^2\) Now use the formula: \[a^2 - b^2 =(a+b)(a-b)\] Here a = x and b = 11..

OpenStudy (anonymous):

What did you get by using that formula ??

OpenStudy (anonymous):

Ok, I get it now! I got: (x + 11)(x - 11) x^(2) - 11x + 11x - 121 x^(2) - 121

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

\[\large (t^3 + 2t^2) - (9t - 18)\] factor out \(t^2\) from first grouping and -9 from second grouping: \[\implies t^2(t + 2) - 9(t + 2)\] \[(t+2)(t^2-9)\]

OpenStudy (anonymous):

Now do the same for \(t^2 - 9\) Factorize it..

OpenStudy (anonymous):

t^(2) - 9 = ... t^(2) - 3^(2)?

OpenStudy (anonymous):

Yes go ahead and have faith you are doing right..

OpenStudy (anonymous):

Thanks again :D

OpenStudy (anonymous):

Welcome..